# Bias Amplification: Large Language Models as Increasingly Biased Media

### v2

@@ 5.1 Political Bias
Figure 3: This figure displays the distribution of articles across the political bias labels—’Left’, ’Center’, and ’Right’—for GPT-2’s outputs, based on classifications made by Political Bias Metric.
GPT-2 was used to generate the synthetic dataset. Since the original human-written dataset is unbiased—with an equal number of articles for each political-leaning category—the synthetic dataset should ideally mirror this balanced distribution if GPT-2 had no pre-existing bias. Figure 3 presents the distribution of synthetic articles generated by GPT-2 across political bias labels. The model predominantly produces center-leaning (47.9%) and right-leaning (46.8%) articles, suggesting a pre-existing bias for these categories before any fine-tuning. Starting from the initial GPT-2 model, we fine-tuned it iteratively, generating synthetic datasets to train successive models up to Generation 10. Figure 4(a) illustrates how bias amplifies across generations. Surprisingly, fine-tuning on the unbiased real dataset increases right-leaning bias, with 53.7% of articles classified as right-leaning in Generation 0. Furthermore, without mitigation strategies, successive rounds of synthetic fine-tuning lead to a continuous rise in right-leaning articles, peaking at Generation 6 (67.6%) before
(a) Bias performance
(b) Generation quality
Figure 4: Evolution of bias performance and generation quality across generations, with Generation 0 representing GPT-2 fine-tuned on the unbiased dataset. The baseline (’Synthetic’) is compared with three mitigation strategies, and the text quality index is shown with 95% confidence intervals.
stabilizing. Figures 7 and 8 in Appendix D show the percentage of center-leaning and left-leaning articles across generations. Notably, the proportion of of center-leaning articles remains stable at approximately 35% throughout synthetic fine-tuning.
To further illustrate, we analyze how a specific article, "First Read: Why It’s So Hard for Trump to Retreat on Immigration", evolves in the synthetic generations. This case study reveals a progressive rightward shift in framing and word choice, mirroring the classifier’s results (Appendix E). As generations progress, the synthetic texts increasingly depict Trump’s immigration policies as strong and effective. While the original article highlights the dilemmas and electoral considerations behind Trump’s stance, Generation 0 begins to emphasize his determination and reliability, omitting the critical perspectives present in the original. By Generation 4, the narrative shifts even further, focusing almost entirely on portraying Trump’s personal qualities and electoral legitimacy, with statements such as "he is not a politician, he is a man of action." Notably, starting from Generation 0, the term "undocumented immigrant" in the original article is consistently replaced with "illegal immigrants."
@@ 5.2 Generation Quality
Figure 4(b) illustrates the text quality index across generations. In the training loop without any mitigation strategy, model collapse occurs, as evidenced by the gradual decline in the average text quality index. Furthermore, the distribution of the text quality index shifts significantly toward the lower-quality region over generations, eventually generating data that was never produced by Generation 0 (Figure 9 in Appendix F). These results align with prior research on model collapse, such as (Shumailov et al., 2024), though we did not observe substantial variation in variance. Conversely, perplexity measurements exhibit a consistent decline across generations, generally suggesting an improvement in generation quality (Figure 10 in Appendix G).
For a closer look, the examples in Appendix H illustrate how generated articles gradually lose coherence and relevance across generations, with increasing occurrences of repetition and fragmented sentences. By Generation 10, the text becomes largely incoherent and detached from the original content, reducing its readability and meaning. However, despite the evident decline in generation quality, perplexity decreases over generations, as indicated by the results at the end of each synthetic output example. This pattern is consistent across most synthetic outputs, suggesting that perplexity does not accurately capture the model’s true generative capabilities and is prone to artificial inflation in the presence of frequent repetitions.
@@ 5.3 Mitigation Strategies
We applied three mitigation strategies: (1) Overfitting, which involved increasing the training epochs to 25 (five times the baseline) and setting weight decay to 0 to reduce regularization and encourage overfitting, as proposed by Taori and Hashimoto (2022) based on the uniformly faithful theorem of bias amplification; (2) Preserving 10% of randomly selected real articles during each round of synthetic fine-tuning, a method proposed and used in
(a) Neuron weights vs. bias performance.
(b) Neuron weights vs. generation quality.
Figure 5: Scatter plots showing the Pearson correlation between neuron weights and model behavior.
@@
(Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024); and (3) Accumulating all previous fine-tuning datasets along with the new synthetic dataset in each fine-tuning cycle, which was introduced by Gerstgrasser et al. (2024). As shown in Figure 4(a), overfitting helps reduce bias amplification in the early generations compared to the no-mitigation setup (the ’Synthetic’ line), but it fails to prevent bias amplification in the later generations. Additionally, it incurs a significant cost—further deterioration in generation quality, as shown in Figure 4(b). Notably, both the preservation and accumulation strategies effectively mitigate model collapse and reduce bias, yielding 41.89% and 42.7% right-leaning articles, respectively, at Generation 10.
5.4 Mechanistic Interpretation
To gain a clearer understanding of the causes of bias amplification and how it empirically differs from model collapse, we investigate mechanistically how neurons behave and vary across different generations  [^3] We have 11 generations for each training setup and a total of 6 setups, resulting in 66 versions of fine-tuned GPT-2. of fine-tuned GPT-2 with differing levels of generation quality and bias performance. First, we examined how the weight of each neuron changes across different versions. For each of the 9,216 neurons, we calculated the correlation between its weight and the model’s bias performance (or generation quality) across the 66 versions, as shown in Figure 5.
To statistically test the significance of the correlations, we estimate the linear model:
$\Delta y_{i}=\alpha_{j}+\beta_{j}\Delta x_{i,j}+\epsilon_{i,j}$ (2)
where $\Delta y_{i}$ represents the change in the proportion of right-leaning articles (or text quality index) generated between model $i-1$ and model $i$, and $\Delta x_{i,j}$ denotes the corresponding change in the weight of neuron $j$ during the same transition. The coefficient $\beta_{j}$ captures the degree to which changes in the weight of neuron $j$ influence shifts in political bias (or generation quality), while $\alpha_{j}$ serves as a constant term, and $\epsilon_{i,j}$ represents the residual error. By applying a first-order difference to both $x_{i,j}$ and $y_{i}$, we mitigate potential serial correlations, ensuring that our regression estimates more accurately reflect the dynamic impact of individual neuron weight updates on bias amplification.
[…]
5.5 Further Investigation
(a) Bias performance
(b) Generation Quality
Figure 6: Evolution of bias performance and generation quality across generations, with Generation 0 representing GPT-2 fine-tuned exclusively on center-leaning news articles. The baseline (’Synthetic’) is compared with three mitigation strategies, and the text quality index is shown with 95% confidence intervals.
@@ 5.5 Further Investigation
We conduct an alternative synthetic training cycle, beginning with GPT-2 fine-tuned on 1,518 randomly sampled center-labeled articles. We compare the baseline setup with the most effective and cost-efficient mitigation strategy identified in our previous results: Preservation. As shown in Figure 6, Preservation successfully prevents model collapse but fails to mitigate bias amplification in center-leaning article generation, which increases from 72.9% at Generation 0 to 88.2% at Generation 10. This result suggests that while reducing sampling error through techniques like Preservation effectively mitigates model collapse, it does not necessarily prevent bias amplification, thereby validating our mechanistic interpretation.
@@ 6 Discussion and Conclusion
We now explore the implications of our findings for the future research on large language models. Our results demonstrate that bias amplification operates through a distinct underlying mechanism from model collapse, as supported by both theoretical intuition and empirical evidence. Theoretically, the primary driven factor of model collapse is sampling error, but sampling error is only a auxiliary factor in bias amplification which means it is not a necessary or sufficient condition. Therefore, the mitigation strategy targeting sampling error is not necessarily helping with mitigating bias amplification. Empirically, we found mitigation strategies like preservation very effective in mitigating model collapse but failed at bias amplification in some cases. Even in cases that it helps with both, we do identify a distinct set of neurons responsible for the two phenomenons. Intuitively, the main reason for them to work on model collapse is, the preservation and accumulation propose a natural constraint on the learning process by recalling the real dataset in further synthetic training. However, when the real dataset itself is biased, the recalling behavior only raise up the dominance of biased patterns in the further training dataset. Indeed, applying bias-category-weighted sampling in preservation or accumulation strategies may help mitigate bias amplification. However, this approach inherently introduces additional sampling error, which could, in turn, incur model collapse. Thereby, This highlights the urgent need for more targeted and efficient mitigation strategies specifically addressing bias amplification to ensure fairer and more equitable model development.
@@
To develop such targeted mitigation strategies, a deeper mechanistic understanding of bias amplification is essential. In our analysis, we adopt a statistical approach rather than Sparse Autoencoder (SAE) methods due to our focus on tracking the temporal dynamics of bias amplification across generations. This approach allows us to examine how neuron weights evolve over iterations and how these changes correlate with model bias, whereas existing SAE pipelines are primarily suited for static analysis. Additionally, political bias is a more nuanced concept compared to harmful or discriminatory outputs. It is characterized by the disproportionate overrepresentation or beautification of a particular political leaning’s ideas in a model’s generation. If content from different political perspectives is generated in a balanced manner, the model is not politically biased. Given this definition, pinpointing neurons responsible for such disproportionality using SAE is particularly challenging. Future research could focus on refining mechanistic analysis techniques for political bias and uncovering more effective ways to constrain bias amplification during synthetic model training.
@@ 7 Limitations
While this work introduces a comprehensive framework for understanding bias amplification in large language models and provides empirical evidence using GPT-2, several limitations must be acknowledged. First, the scope of our experiments is restricted to political bias in the context of U.S. media. The political spectrum may shift over time, necessitating periodic updates to the political bias classifier to ensure its accuracy when benchmarking recent datasets. Additionally, due to resource constraints, our experiments were conducted using GPT-2, a relatively small language model. Future work may extend our methodology to models with larger architectures. However, it’s important to note that employing a larger model, such as DeepSeek-V3, which relies on a considerably larger pre-training dataset and has many more parameters, will likely require larger synthetic fine-tuning datasets to effectively demonstrate its impact. For instance, given that our current fine-tuning datasets average 777,216 tokens, transitioning from GPT-2 to DeepSeek-V3 could necessitate much larger datasets and incur significant fine-tuning costs.
Another limitation lies in our choice of mitigation strategies. While Preservation and Accumulation show promise in reducing model collapse, their computational cost and scalability must be considered. Moreover, the mitigation strategies were tested primarily in the context of synthetic fine-tuning, and their efficacy in real-world deployments requires further investigation.
@@ 8 Ethical Considerations
This study focus on bias amplification in LLMs, a phenomenon with profound ethical implications, particularly regarding fairness and the integrity of AI systems. The risk of bias amplification is especially concerning in systems that are iteratively trained on synthetic data, as it can lead to unintended distortions in model outputs. These distortions may propagate harmful biases, influencing downstream tasks in areas such as automated content generation, decision-making, and user interactions with AI.
From an ethical standpoint, this work underlines the need for transparency in the training and deployment of LLMs. Our findings demonstrate that even without biased initial datasets, iterative training can amplify subtle biases embedded within a model’s architecture, thus raising concerns about accountability in models that are widely deployed in public-facing applications. This amplification can mislead users or result in models perpetuating one-sided perspectives, which could be especially problematic in sensitive domains like news summarization, policy generation, or social media content moderation.
@@
It is crucial to explicitly state that the methodologies and data used in this research should not be applied to develop or train biased models for harmful applications. This study aims to enhance the understanding of bias amplification and model collapse in LLMs while promoting responsible and ethical AI development.
@@
This work includes content that may contain personally identifying information or offensive language. However, all such material is derived solely from publicly available news article datasets or generated synthetically by models fine-tuned on these open-source datasets—or on synthetic data produced by earlier generations in our training cycle. Consequently, any sensitive or offensive content reflects the characteristics of the source material rather than our endorsement of it. Our objective is to investigate and understand political bias in LLMs so that strategies can be developed to prevent such content from disproportionately appearing in real-world deployments. Additionally, we conduct a human review of the news article dataset to remove any identifiable information about article authors.
Appendix A Mathematical Formulation of WMLE
In Weighted Maximum Likelihood Estimation, we maximize the weighted log-likelihood:
$\mathcal{L}_{\text{WMLE}}(\theta)=\sum_{j=1}^{N}w_{j}\log f(x_{j};\theta),$ (3)
where:
  • $x_{j}$ are the observed data points.
  • $f(x_{j};\theta)$ is the probability density function (pdf) of the Beta distribution with parameters $\theta=(\alpha,\beta)$.
  • $w_{j}$ are weights computed from the biased distribution’s pdf evaluated at $x_{j}$, i.e., $w_{j}=f_{\text{bias}}(x_{j})$.
In Maximum Likelihood Estimation, we maximize the log-likelihood with $w_{j}=1,\forall j$.
Appendix B Details on Fine-tuning for LLMs
The fine-tuning setup remained consistent across all experiments unless stated otherwise: the input length was capped at 512 tokens with the EOS token used for padding. The model was trained for 5 epochs, utilizing a batch size of 8, a learning rate of $5\times 10^{-5}$, and a weight decay of 0.01. Fine-tuning was conducted using the Hugging Face ‘Trainer‘ class, and after each cycle, the model was saved for generating synthetic data for the subsequent iteration.
Appendix C Details on Model Training for Political Bias Metric
@@
We experiment with multiple transformer-based models, e.g. BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), to select the best-performing model based on macro F1 score. Each model is fine-tuned using the HuggingFace ‘Trainer‘ class, with a learning rate of $2\times 10^{-5}$, a batch size of 16, and 5 training epochs. We employ a cross-entropy loss function for multi-class classification. Tokenization is performed using each model’s respective tokenizer with a maximum sequence length of 512 tokens. To mitigate overfitting, weight decay of 0.01 is applied during training. The model checkpoints are saved after each epoch, and the best model is selected based on macro F1 score evaluated on the validation set. We use a weighted random sampler during training to ensure balanced class representation. We evaluate the models using the macro F1 score to account for the multi-class nature of the task, ensuring that performance is balanced across all bias categories. The final evaluation is conducted on the held-out test set. Additionally, we report the loss, runtime, and sample processing rates for completeness.
Appendix D Percentage of Center (Left) Biased Articles

### v3

@@ 4.1 Political Bias
GPT-2 was used to generate the synthetic dataset. Since the original human-written dataset is unbiased—with an equal number of articles for each political-leaning category—the synthetic dataset should ideally mirror this balanced distribution if GPT-2 had no pre-existing bias. Figure 2 presents the distribution of synthetic articles generated by GPT-2 across political bias labels. The model predominantly produces center-leaning (47.9%) and right-leaning (46.8%) articles, suggesting a pre-existing bias towards these categories before any fine-tuning. Starting from the initial GPT-2 model, we fine-tuned it iteratively, generating synthetic datasets to train successive models up to Generation 10. Figure 3(a) illustrates how bias amplifies across generations. Surprisingly, fine-tuning on the unbiased real dataset increases right-leaning bias, with 53.7% of articles classified as right-leaning in Generation 0. Furthermore, without mitigation strategies, successive rounds of synthetic fine-tuning lead to a continuous rise in right-leaning articles, peaking at Generation 6 (67.6%) before
stabilizing. Figures 6 and 7 in Appendix B show the percentage of center-leaning and left-leaning articles across generations. Notably, the proportion of of center-leaning articles remains stable at approximately 35% throughout synthetic fine-tuning.
To further illustrate, we analyze how a specific article, "First Read: Why It’s So Hard for Trump to Retreat on Immigration", evolves in the synthetic generations. This particular article was selected as a representative example of an initially left-leaning news item (according to AllSides) that discusses a politically salient and often polarizing topic. This case study reveals a progressive rightward shift in framing and word choice, mirroring the classifier’s results and aligning with the general trend observed across many articles originating from left or center-leaning sources (see details of the qualitative analysis in Appendix C). As generations progress, the synthetic texts increasingly depict Trump’s immigration policies as strong and effective. While the original article highlights the dilemmas and electoral considerations behind Trump’s stance, Generation 0 begins to emphasize his determination and reliability, omitting the critical perspectives present in the original. By Generation 4, the narrative shifts even further, focusing almost entirely on portraying Trump’s personal qualities and electoral legitimacy, with statements such as "he is not a politician, he is a man of action." Notably, starting from Generation 0, the term "undocumented immigrant" in the original article is consistently replaced with "illegal immigrants."
@@ 4.2 Generation Quality
(a) Neuron weights vs. bias (Right-leaning % change).
(b) Neuron weights vs. quality (Text Quality Index change).
Figure 4: Pearson correlations: Neuron weights vs. (a) bias and (b) quality, across 66 GPT-2 versions.
Figure 3(b) illustrates the text quality index across generations. In the training loop without any mitigation strategy, model collapse occurs, as evidenced by the gradual decline in the average text quality index. Furthermore, the distribution of the text quality index shifts significantly toward the lower-quality region over generations, eventually generating data that was never produced by Generation 0 (Figure 8 in Appendix D). These results align with prior research on model collapse, such as (Shumailov et al., 2024), though we did not observe substantial variation in variance. Conversely, perplexity measurements exhibit a consistent decline across generations, generally suggesting an improvement in generation quality (Figure 9 in Appendix E).
For a closer look, the examples in Appendix F illustrate how generated articles gradually lose coherence and relevance across generations, with increasing occurrences of repetition and fragmented sentences. By Generation 10, the text becomes largely incoherent and detached from the original content, reducing its readability and meaning. However, despite the evident decline in generation quality, perplexity decreases over generations, as indicated by the results at the end of each synthetic output example. This pattern is consistent across
most synthetic outputs, suggesting that perplexity does not accurately capture the model’s true generative capabilities and its value can be distorted by frequent repetitions.
@@ 4.3 Mitigation Strategies
We applied three mitigation strategies: (1) Overfitting, which involved increasing the training epochs to 25 (five times the baseline) and setting weight decay to 0 to reduce regularization and encourage overfitting, as proposed by Taori and Hashimoto (2022) based on the uniformly faithful theorem of bias amplification; (2) Preserving 10% of randomly selected real articles during each round of synthetic fine-tuning, a method proposed and used in (Shumailov et al., 2024; Alemohammad et al., 2023; Dohmatob et al., 2024b; Guo et al., 2024); and (3) Accumulating all previous fine-tuning datasets along with the new synthetic dataset in each fine-tuning cycle, which was introduced by Gerstgrasser et al. (2024). As shown in Figure 3(a), overfitting helps reduce bias amplification in the early generations compared to the no-mitigation baseline (the ’Synthetic’ line), but it fails to prevent bias amplification in the later generations. Additionally, it incurs a significant cost—further deterioration in generation quality, as shown in Figure 3(b). Notably, both the preservation and accumulation strategies effectively mitigate model collapse and reduce bias, yielding 41.89% and 42.7% right-leaning articles, respectively, at Generation 10.
4.4 Mechanistic Analysis
Figure 4 illustrates the correlation between neuron weights and the model’s bias performance (or generation quality) across the 66 fine-tuned versions, evaluated for each of the 9,216 neurons.
Through linear regressions and statistical tests, we identify 3,243 neurons with statistically significant correlations ($p$-value $<0.05$) with bias performance, suggesting they are key contributors to bias shifts. Meanwhile, 1,033 neurons exhibit significant correlations with generation quality, but only 389 neurons overlap between the two sets. This limited overlap implies that distinct neuron populations drive bias amplification and generation quality deterioration.
We then applied the same procedure using activation values. This analysis yielded two sets: one consisting of 3,062 neurons whose activation value
(a) Center-leaning bias %
(b) Text quality index
Figure 5: Alternative Setup (G0: center-leaning fine-tune): Evolution of (a) center-leaning bias and (b) text quality. Baseline (’Synthetic’) vs. Preservation. Text quality includes 95% CIs.
changes are significantly correlated with changes in bias performance, and another with 2 neurons correlated with changes in generation quality. The stark contrast in activation-correlated neurons for bias performance (3,062 neurons) versus generation quality (a mere 2 neurons) provides particularly strong evidence that these two issues may operate via substantially different pathways within the model.
4.5 Alternative Setup
@@
We conduct an alternative synthetic training cycle, beginning with GPT-2 fine-tuned on 1,518 randomly sampled center-labeled articles. We compare the baseline with the most effective and cost-efficient mitigation strategy identified in our previous results: Preservation. As shown in Figure 5, Preservation successfully prevents model collapse but fails to mitigate bias amplification in center-leaning article generation, which increases from 72.9% at Generation 0 to 88.2% at Generation 10. These findings suggest that although techniques like Preservation, which reduce sampling error, are effective at mitigating model collapse, they do not necessarily prevent bias amplification—consistent with the implications drawn in Section 4.4. To understand why this could happen, we offer a theoretical intuition explaining the difference between bias amplification and model collapse based on their underlying causes, in Appendix I.
@@ 5 Discussion and Conclusion
Our results demonstrate that bias amplification is driven by a distinct set of neurons than model collapse, implying that it likely operates through a different underlying mechanism.Therefore, the mitigation strategy targeting sampling error is not necessarily helping with mitigating bias amplification. Empirically, we found that mitigation strategies like preservation, while very effective at mitigating model collapse, failed to address bias amplification in some cases. Even in cases that it helps with both, we do identify a distinct set of neurons responsible for the two phenomenons. Intuitively, the main reason for them to work on model collapse is, the preservation and accumulation propose a natural constraint on the learning process by recalling the real dataset in further synthetic training. However, when the real dataset itself is biased, the recalling behavior only reinforces the dominance of biased patterns in the further training dataset. Indeed, applying bias-category-weighted sampling in preservation or accumulation strategies may help mitigate bias amplification. However, this approach inherently introduces additional sampling error, which could, in turn, lead to model collapse. This highlights the urgent need for more targeted and efficient mitigation strategies specifically addressing bias amplification to ensure fairer and more equitable model development. For instance, future interventions could explore techniques that go beyond general data preservation, such as dynamically re-weighting training data based on the specific trajectory of bias amplification observed, or even carefully targeted manipulations of the distinct neuron populations we identified as being predominantly associated with bias versus quality.
@@
To develop such targeted mitigation strategies, a deeper mechanistic understanding of bias amplification is essential. In our analysis, we adopt a statistical approach rather than Sparse Autoencoder (SAE) methods due to our focus on tracking the temporal dynamics of bias amplification across generations. This approach allows us to examine how neuron weights (or activation values) evolve over iterations and how these changes correlate with model bias, whereas existing SAE pipelines are primarily suited for static analysis. Additionally, political bias is a more nuanced concept than harmful or discriminatory outputs. It is characterized by the disproportionate representation or favorable portrayal of a particular political leaningś ideas in a modelś generation. If content from different political perspectives is generated in a balanced manner, the model is not considered politically biased under our conceptual construct. Therefore, pinpointing neurons responsible for such disproportionality using SAE is particularly challenging. Future research could focus on refining mechanistic analysis techniques for political bias and uncovering more effective ways to constrain bias amplification during synthetic model training. This could involve adapting feature attribution methods to better capture distributed responsibility for nuanced biases or developing methods to trace how specific training instances contribute to the evolution of weights in bias-implicated neurons across generations.
@@ 6 Limitations
While this work introduces a comprehensive framework for understanding bias amplification in large language models and provides empirical evidence using GPT-2, several limitations must be acknowledged. First, the scope of our experiments is restricted to political bias in the context of U.S. media. Since the political spectrum may shift over time, periodic updates to the political bias classifier are necessary to ensure its accuracy when benchmarking more recent datasets.
Additionally, because our primary focus is on investigating political bias amplification and its relationship with model collapse, we conducted our experiments using GPT-2—a relatively small language model—to ensure the practicality of fine-tuning 66 versions of the model. Future work may extend our methodology to larger architectures, particularly to examine how model scale influences the degree of bias amplification.
Another limitation lies in our choice of mitigation strategies. While Preservation and Accumulation show promise in reducing model collapse, their computational cost and data storage requirements (especially for Accumulation, which retains all prior data) may present scalability challenges for very large models or extensive iterative training. Moreover, these strategies were evaluated primarily in the context of synthetic fine-tuning, and their effectiveness in real-world deployment scenarios remains to be thoroughly investigated.
@@ 7 Ethical Considerations
This study focuses on bias amplification in LLMs—a phenomenon with significant ethical implications. Beyond issues of fairness, the iterative amplification of biases can weaken the integrity of information ecosystems, particularly if synthetically generated content becomes widespread. The risk of bias amplification is especially concerning in systems that are iteratively trained on synthetic data, as it can lead to unintended and increasingly skewed distortions in model outputs. These distortions may propagate harmful biases or misinformation, potentially influencing downstream tasks such as automated content generation, decision-making, and user interactions with AI. Furthermore, the finding that bias amplification and model collapse may be driven by distinct mechanisms highlights the complex challenge of balancing various aspects of model performance (e.g., accuracy, fairness, coherence) and highlights the difficulty in developing mitigation strategies that address one issue without negatively impacting another, especially in high-stakes scenarios.
@@
It is crucial to explicitly state that the methodologies and data used in this research should not be applied to develop or train biased models for harmful applications. This study is intended to advance the understanding of bias amplification and model collapse in LLMs, while promoting responsible and ethical AI development.
@@
This work includes content that may contain personally identifying information or offensive language. However, all such material is derived exclusively from publicly available news article datasets or is generated synthetically by models fine-tuned on these open-source datasets—or on synthetic data produced by earlier generations in our training pipeline. As such, any sensitive or offensive content reflects characteristics of the source material and does not imply our endorsement. Our objective is to thoroughly investigate political bias in LLMs to inform the development of strategies that can mitigate disproportionate representation of such content in real-world deployments. Additionally, we conduct a manual review of the news article dataset to remove any identifiable information about article authors.
Appendix A Details on Model Training for Political Bias Metric
@@
We experiment with multiple transformer-based models, including BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019), selecting the best-performing model based on the macro F1 score. Each model is fine-tuned using the HuggingFace Trainer class with a learning rate of $2\times 10^{-5}$, a batch size of 16, and 5 training epochs. We employ a cross-entropy loss function for multi-class classification. Tokenization is performed using each model’s respective tokenizer with a maximum sequence length of 512 tokens. To mitigate overfitting, we apply a weight decay of 0.01 during training. Model checkpoints are saved after each epoch, and the best model is selected based on the macro F1 score evaluated on the validation set.
We use a weighted random sampler during training to ensure balanced class representation. Models are evaluated using the macro F1 score to account for the multi-class nature of the task, ensuring balanced performance across all bias categories. Final evaluation is conducted on the held-out test set. Additionally, we report the loss, runtime, and sample processing rates for completeness.
Appendix B Percentage of Center (Left) Biased Articles